OpenStudy (anonymous):

Solve . Answer x = 5 x = -5 No Solution x = 3

OpenStudy (anonymous):



OpenStudy (anonymous):

@ghazi

OpenStudy (anonymous):

x = 1328/5

OpenStudy (ghazi):

\[\sqrt[3]{5x+2}-3=8\] then \[\sqrt[3]{5x+2}=8+3=11\] cube both the sides

OpenStudy (anonymous):

cuberoot(5x+2) =3 cubing both sides we have 5x+2 = 3^3 or 5x+2=27 or 5x=27-2 or 5x=25 or x=5

OpenStudy (anonymous):

5

OpenStudy (ghazi):

\[5x+2=1331\] , you're gonna get x=5

OpenStudy (ghazi):

there is a mistake in my last comment